OpenStudy (anonymous):

IF y=-5 when x=15, find y when x=-1. PLEASE HELP ME

Directrix (directrix):

Is this a direct variation problem?

OpenStudy (anonymous):

YES

Directrix (directrix):

The form for that is y = (1/k) x and you first have to find the variation constant k. So, -5 = (1/k) (15) and k = -3.

Directrix (directrix):

Using the equation y = ( -1/3) (-1), then y = 1/3 when x = -1/3.